# Privacy Preserving Mechanisms for Coordinating Airspace Usage in Advanced Air Mobility

### v1

@@ 4.1 Step 1: Distributed Algorithm for Computing Fractional-Competitive Equilibrium
To compute the fractional-competitive equilibrium, we propose Algorithm 1 to solve the fixed-point problem described in Proposition 3.5 in a distributed manner. Algorithm 1 emulates the fixed-point iteration for the mapping
@@
Since the SP does not have access to $\lambda^{\dagger}(\omega)$ as it does not know the private valuation of AAM vehicles, we employ a two-loop approach. In the inner loop, we iteratively solve the convex optimization problem (3) in a distributed manner that does not require the SP to access the private valuations of AAM vehicles. This is done by repeatedly interacting with the AAM vehicles for finite number of rounds, while holding $\omega$ constant, to approximate $\lambda^{\dagger}(\omega)$. This approximation is then used to update $\omega$ in the outer loop.
To solve the inner loop problem in a distributed manner, we reformulate (3) as (4) by introducing two additional variables, ${\mathbf{y}}$ and ${\mathbf{z}}$. This reformulation enables the use of distributed optimization techniques, facilitating distributed computation across multiple agents while preserving the structure of the original problem.
@@
$\quad\sum_{u\in U}y_{u,e}+\bar{z}_{e}=\ell_{e}\quad\forall\ e\in\tilde{\mathcal{E}}$ (4c)
@@
In this reformulation, Equation (4b) enforces equality between ${\mathbf{x}}$ and ${\mathbf{y}}$, while Equation (4c) ensures that capacity constraints are met. Constraints (4d)-(4f) are identical to (3c)-(3e) along with additional positivity constraints on ${\mathbf{y}},{\mathbf{z}}.$ This reformulation ensures that the Lagrangian of (4) becomes a separable function of $\bar{\mathbf{x}}_{u}$, enabling the problem to be solved in a distributed manner using the ADMM algorithm (He et al., 2023; Jalota et al., 2023). The variable ${\mathbf{y}}$ can be interpreted as the ”baseline demand” expected by the service provider (SP), while ${\mathbf{z}}$ represents the ”resource surplus” in each region. Next, we describe the inner and outer loop in detail. We index the inner loop iterations by $n=1,2,...,N$ and the outer loop iterations by $k=1,2,...,K$.
Inner Loop: Given that outer loop is at iteration $k$, at any iteration $n$ of the inner loop: $(a)$ every AAM vehicle $u$ keeps track its individual demand $\bar{\mathbf{x}}^{(n,k)}_{u}$, and $(b)$ the SP keeps track of three quantities: an estimate of expected baseline demands ${\mathbf{y}}^{(n,k)}$, expected resource surplus ${\mathbf{z}}^{(n,k)}$, price of all regions $\mathbf{p}^{(n,k)}$, and a dual multiplier $\lambda^{(n,k)}$ which is used to adjust budgets of AAM vehicles.
Local update for each AAM vehicle: Given that the outer loop is at iteration $k$, at every iteration $n$ of the inner loop, each AAM vehicle $u$ receives its expected baseline demand ${\mathbf{y}}^{(n,k)}_{u}$, the current prices on regions $\mathbf{p}^{(n,k)}$ and the dual multiplier corresponding to its local constraints $\lambda^{(n,k)}_{u}$. Using this information, AAM vehicle updates its requested demand using the following update and shares this with SP:
[…]
$\quad\hskip 56.9055pt-\frac{\beta}{2}(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}-1)^{2}-\frac{\beta}{2}\|\mathbf{y}_{u}^{(n,k)}-\mathbf{x}_{u}\|^{2}\Bigg{)}.$
Remark 4.1.
The update in Equation (5) can be interpreted as a “proxy bidding agent” used by each AAM vehicle. This proxy attempts to maximize the vehicle’s budget-adjusted utility while penalizing deviations from the baseline demand, overspending artificial currency, and violating constraints. These constraints capture the requirement that the bundle of edges selected by the AAM vehicle results in a feasible path on the time-extended graph.
Updates by SP: Using the demand from AAM vehicles, the SP updates the baseline demand and the excess supply as follows:
$(\mathbf{y}^{(n+1,k)},{\mathbf{z}}^{(n+1,k)})=\operatorname*{arg\,max}_{{\mathbf{z}}\geq\mathbf{0},{\mathbf{y}}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}}\Bigg{(}-\frac{\beta}{2}\|\mathbf{y}_{u}-\mathbf{x}_{u}^{(n+1,k)}\|^{2}-\frac{\beta}{2}\|\sum_{u\in U}\mathbf{y}_{u}+\mathbf{z}-\ell\|^{2}-\sum_{e\in\tilde{\mathcal{E}}}p^{(n,k)}_{e}{z}_{e}\Bigg{)}.$ (6)
The objective in (6) requires that SP minimizes three terms: $(i)$ difference between expected baseline demand by SP and the demand sent by AAM vehicles; $(ii)$ constraint violation at the resources; and $(iii)$ unused capacity is minimized on any resource with a positive price. Furthermore, the SP updates the price estimates using the updated values of baseline demand and excess supply as follows:
The above equations resemble the idea that an SP should increase the price if the capacity constraint is violated and reduce it if there is available capacity. Finally, the SP updates the dual multiplier estimate $\lambda_{u},$ for every AAM vehicle $u$ as follows:
$\lambda_{u}^{(n+1,k)}={\lambda}_{u}^{(n,k)}+\beta\left(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}^{(n+1,k)}+x_{u,\varnothing}^{(n+1,k)}-1\right),$ (8)
which keeps track of flow constraint violations by every agent.
Outer loop. In the outer loop, we update the budget adjustment after every $N$ iteration of the inner loop, using the value of $\lambda^{(N,k)}$ to approximate the fixed-point iteration (line 12 of Algorithm 1). This step ensures that budget adjustments progressively converge toward the equilibrium by iteratively refining the dual variables based on the current solution from the inner loop.
Termination criterion: We terminate the algorithm once all of the following errors fall below a predefined threshold:
  • Market clearing error (MCE): Smaller values of MCE ensure that resources with a positive price maintain a balance between demand and supply, while resources priced at zero satisfy capacity constraints. We define
    $\textsf{MCE}(\bar{\mathbf{x}},\mathbf{p})=\sqrt{\sum_{e\in\tilde{\mathcal{E}}:p_{e}>0}\left({z}_{e}\right)^{2}+\sum_{e\in\tilde{\mathcal{E}}:p_{e}=0}\left(\max\left\{0,{z}_{e}\right\}\right)^{2}},$ (9)
    where ${z}_{e}=\sum_{u\in U}x_{u,e}-\ell_{e}$ is the over demand.
    $\textsf{ICE}(\bar{\mathbf{x}})=\sqrt{\sum_{u\in U}(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}-1)^{2}}.$ (10)
  • Baseline demand error (BDE): Smaller value of BDE ensure that (4b) is satisfied. We define
    $\textsf{BDE}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{u\in U}\|\mathbf{y}_{u}-\mathbf{x}_{u}\|_{2}^{2}}.$ (11)
@@
2:for $k=1,2,...,K$ do
3:    for $n=0,1,....N-1$ do
@@
4:         Upon observing $\mathbf{p}^{(n,k)},$ $\mathbf{y}^{(n,k)},$$\omega_{u}^{(k)}$, each AAM vehicle independently updates its demand using (5)
5:         Upon observing $\bar{\mathbf{x}}_{u}^{(n,k)}$, SP updates the baseline demand and overdemand using (6)
6:         SP updates the price estimates using (7)
7:         SP tracks the dual multiplier estimate corresponding to $\omega^{(k)}$ using (8)
8:         if $\max\{\textsf{MCE}(\bar{\mathbf{x}}^{(n,k)},{\mathbf{p}}^{(n,k)}),\textsf{ICE}(\bar{\mathbf{x}}^{(n,k)}),\textsf{BDE}(\mathbf{x}^{(n,k)},\mathbf{y}^{(n,k)})\}\leq\textsf{tol}$ then
9:             Return($\bar{\mathbf{x}}^{(n,k)},{\mathbf{p}}^{(n,k)}$)
10:         end if
11:    end for
12:    $\omega^{(k+1)}=\lambda^{(N,k)}$
13:    $\mathbf{p}^{(0,k+1)}\leftarrow\mathbf{p}^{(N,k)},\mathbf{y}^{(0,k+1)}\leftarrow\mathbf{y}^{(N,k)}$
14:end for
@@ 4.2 Step 2: Computing Integral Allocation
Using the fractional allocations from Algorithm 1, the service provider (SP) computes an integral allocation in a distributed manner using Algorithm 2. The SP sets the airspace price to $\mathbf{p}^{\dagger}$ (the output of Algorithm 1) and generates a ranked list of AAM vehicles based on $\bar{\mathbf{x}}^{\dagger}$. This list is created by ranking the AAM vehicles in descending according to the value of the (fractional) allocation of their preferred routes.
Ranked List: More formally, define $s^{\ast}(u)$ to denote the most desired route of AAM vehicle $u$ in $R_{u}$. The SP creates a ranking over agents based on decreasing values of $x^{\dagger}_{u,e^{\ast}(s^{\ast}(u))}$, where $e^{\ast}(s^{\ast}(u))$ denotes the departing edge on the route $s^{\ast}(u)$. We denote the ranked list777ties are broken arbitrarily. of AAM vehicles by rank.
Integral Allocation: The SP goes over each AAM vehicle as per their order in rank and retrieves the most desired route for each vehicle (determined by solving the individual optimization problem (2)). If any capacity constraints are violated at a resource along the route, the SP either removes that resource from the available pool for all remaining agents or, alternatively, increases the price of that resource to infinity for the remaining AAM vehicles (as described in line 13 of Algorithm 2).
Algorithm 2 Ranking System
@@
2:for $i=1$ to $|U|$ do
3:    $u\leftarrow\textsf{{rank}}_{i}$
@@
6:         AAM vehicle $u$ reports $\bar{\mathbf{x}}_{u}$ by solving (2)
7:         if $x_{u,e}\leq\ell_{e}$ for every $e\in\tilde{\mathcal{E}}$ then
8:             Update available capacity of regions: $\ell\leftarrow\ell-\mathbf{x}_{u}$
@@
10:         else
11:             Contested_goods = $e\in\tilde{\mathcal{E}}$$x_{u,e}>\ell_{e}$}

### v2

@@ 5.1 Step 1: Distributed Algorithm for Computing Fractional-Competitive Equilibrium
To compute a fractional-competitive equilibrium, we propose Algorithm 1 to solve the fixed-point problem described in Proposition 4.5 in a distributed manner. Algorithm 1 emulates the fixed-point iteration for the mapping
@@
Since the SP lacks access to $\lambda^{\dagger}(\omega),$ as computing it requires solving (3), which in turn depends on the private valuations of AAM vehicles, we adopt a two-loop approach to circumvent this challenge. In the inner loop, we iteratively solve the convex optimization problem (3) in a distributed manner that does not require the SP to access private valuations of AAM vehicles. This is done by repeatedly interacting with the AAM vehicles for a finite number of rounds, while holding $\omega$ constant, to approximate $\lambda^{\dagger}(\omega)$. This approximation is then used to update $\omega$ using a fixed point iteration in the outer loop.
To solve the inner-loop problem in a distributed manner, we reformulate (3) into (4) by introducing two additional variables, ${\mathbf{y}}$ and ${\mathbf{z}}$. This reformulation enables the use of distributed optimization techniques, facilitating distributed computation across multiple agents while preserving the structure of the original problem.
@@
$\quad\sum_{u\in U}y_{u,e}+z_{e}=\ell_{e}\quad\forall\ e\in\tilde{\mathcal{E}}$ (4c)
@@
In this reformulation, Equation (4b) enforces the equality between ${\mathbf{x}}$ and ${\mathbf{y}}$, while Equation (4c) ensures that capacity constraints are met. Constraints (4d)-(4f) is identical to (3c)-(3e) along with additional positivity constraints on ${\mathbf{y}},{\mathbf{z}}.$ This reformulation ensures that the Lagrangian of (4) becomes a separable function of $\bar{\mathbf{x}}_{u}$, allowing the problem to be solved in a distributed manner using the ADMM algorithm (He et al., 2023; Jalota et al., 2023). The variable ${\mathbf{y}}$ can be interpreted as the “expected allocation” estimate of the service provider (SP), while ${\mathbf{z}}$ represents the “resource surplus” in each region. Next, we describe the inner and outer loops in detail. We index the inner loop iterations by $n=1,2,...,N$ and the outer loop iterations by $k=1,2,...,K$. A flowchart of our algorithm in Step 1 is shown in Figure 4.
Figure 4: Flowchart of Algorithm 1, illustrating the processes executed independently by the SP and AAM vehicles, as well as the steps computed within the inner and outer loops.
@@
2:for $k=0,1,2,...,K-1$ $\triangleright$ Outer loop starts
3:    for $n=0,1,...,N-1$ $\triangleright$ Inner loop starts
4:Distributed Updates:
@@
5:         Using $\mathbf{p}^{(n,k)},$ $\mathbf{y}_{u}^{(n,k)},\lambda_{u}^{(n,k)},$ $\omega_{u}^{(k)}$, each AAM vehicle independently updates its demand $\bar{\mathbf{x}}_{u}^{(n+1,k)}$ as follows
[…]
$\quad\hskip 56.9055pt-\frac{\beta}{2}(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}-1)^{2}-\frac{\beta}{2}\|\mathbf{y}_{u}^{(n,k)}-\mathbf{x}_{u}\|^{2}\Bigg{)},$
where $\beta$ is a positive scalar that represents the step-size in the ADMM algorithm (cf. Appendix C).
6:Updates by Service Provider (SP):
7:         Using $(\bar{\mathbf{x}}_{u}^{(n+1,k)})_{u\in U}$, SP updates the expected allocation $\mathbf{y}^{(n+1,k)}$ and resource surplus $\mathbf{z}^{(n+1,k)}$ as follows:
$(\mathbf{y}^{(n+1,k)},{\mathbf{z}}^{(n+1,k)})=\operatorname*{arg\,max}_{{\mathbf{y}}\in\mathbb{R}^{U\times|\tilde{\mathcal{E}}|},\ {\mathbf{z}}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}_{+}}\Bigg{(}-\frac{\beta}{2}\sum_{u\in U}\|\mathbf{y}_{u}-\mathbf{x}_{u}^{(n+1,k)}\|^{2}-\frac{\beta}{2}\|\sum_{u\in U}\mathbf{y}_{u}+\mathbf{z}-\ell\|^{2}-\sum_{e\in\tilde{\mathcal{E}}}p^{(n,k)}_{e}{z}_{e}\Bigg{)}.$ (6)
8:         Using $(\mathbf{y}^{(n+1,k)},\mathbf{z}^{(n+1,k)}),$ SP updates the price estimates as follows
9:         For every $u\in U,$ using $\bar{\mathbf{x}}_{u}^{(n+1,k)},$ SP updates the dual multiplier $\lambda_{u}^{(n+1,k)}$ as follows:
$\lambda_{u}^{(n+1,k)}={\lambda}_{u}^{(n,k)}+\beta\left(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}^{(n+1,k)}+x_{u,\varnothing}^{(n+1,k)}-1\right).$ (8)
10:         if $\textsf{CE}(\bar{\mathbf{x}}^{(n+1,k)},{\mathbf{p}}^{(n+1,k)})\leq\textsf{tol}_{\textsf{CE}},\textsf{ICE}(\bar{\mathbf{x}}^{(n+1,k)})\leq\textsf{tol}_{\textsf{ICE}},\textsf{EAE}(\mathbf{x}^{(n+1,k)},\mathbf{y}^{(n+1,k)})\leq\textsf{tol}_{\textsf{EAE}}$ then
11:             Return $\bar{\mathbf{x}}^{\dagger}=\bar{\mathbf{x}}^{(n+1,k)},\bar{\mathbf{p}}^{\dagger}={\mathbf{p}}^{(n+1,k)}$
12:         end if
13:    end $\triangleright$ Inner loop ends
14:    For every $u\in U,$ SP updates the $\omega^{(k+1)}=\lambda^{(N,k)}$
15:    $\mathbf{p}^{(0,k+1)}\leftarrow\mathbf{p}^{(N,k)},\mathbf{y}^{(0,k+1)}\leftarrow\mathbf{y}^{(N,k)}$
16:end $\triangleright$ Outer loop ends
17:Return $\bar{\mathbf{x}}^{\dagger}=\bar{\mathbf{x}}^{(N,K)},\bar{\mathbf{y}}^{\dagger}={\mathbf{p}}^{(N,K)}$
Inner Loop: Inner loop iterations are obtained by performing ADMM iterations [^8] Derivation of ADMM updates for (4) is provided in Appendix C. for (4) with step-size parameter $\beta$. For any $\omega\in\mathbb{R}_{\geq 0}^{|U|},$ this implementation allows us to estimate the dual multiplier $\lambda^{\dagger}(\omega)$ in a distributed manner without requiring knowledge of the private valuations of AAM vehicles. Next, we describe these iterations in more detail.
Given that outer loop is at iteration $k$, at any iteration $n$ of the inner loop: $(a)$ every AAM vehicle $u$ keeps track of its individual demand $\bar{\mathbf{x}}^{(n,k)}_{u}$, and $(b)$ the SP keeps track of three quantities: an estimate of expected allocations ${\mathbf{y}}^{(n,k)}$, expected resource surplus ${\mathbf{z}}^{(n,k)}$, price of all regions $\mathbf{p}^{(n,k)}$, and a dual multiplier $\lambda^{(n,k)}$ which is used to adjust budgets of AAM vehicles.
Local update for each AAM vehicle: Given that the outer loop is at iteration $k$, at every iteration $n$ of the inner loop, each AAM vehicle $u$ receives its expected expected allocation ${\mathbf{y}}^{(n,k)}_{u}$, the current prices on regions $\mathbf{p}^{(n,k)}$ and the dual multiplier corresponding to its local constraints $\lambda^{(n,k)}_{u}$. Using this information, AAM vehicle updates its requested demand using (5) and shares this with SP.
Remark 5.1.
The update in Equation (5) can be implemented through a “proxy bidding agent” in place of an actual AAM vehicle. The AAM vehicle operator can have the proxy agent at their local end where they feed AAM vehicle valuation and then the proxy agent participates on behalf of AAM vehicle, which ensures that no on-board energy resources are used to run Algorithm 1. This proxy attempts to maximize the vehicle’s budget-adjusted utility while penalizing deviations from the expected allocation, overspending artificial currency, and violating constraints. These constraints capture the requirement that the bundle of edges selected by the AAM vehicle results in a feasible path on the time-extended graph.
Updates by SP: Using the demand from AAM vehicles, the SP updates the expected allocation and the excess supply through (6). The objective in (6) requires that SP minimize three terms: $(i)$ difference between the expected allocation by SP and the demand sent by AAM vehicles; $(ii)$ violation of constraints in the resources; and $(iii)$ the unused capacity is minimized on any resource with a positive price.
Next, the SP updates the price estimates using the updated values of the expected allocation and the excess supply through (7). Equation (7) resembles the idea that an SP should increase the price if the capacity constraint is violated and reduce it if there is available capacity.
Finally, the SP updates the dual multiplier estimate $\lambda_{u},$ for every AAM vehicle $u$ using (8).
Outer loop. In the outer loop, we update the budget adjustment after every $N$ iteration of the inner loop, using the value of $\lambda^{(N,k)}$ to approximate the fixed-point iteration (line 12 of Algorithm 1). This step ensures that budget adjustments progressively converge toward equilibrium by iteratively refining the dual variables based on the current solution from the inner loop.
Termination criterion: We terminate the algorithm once all of the following errors fall below their predefined threshold:
  • Complementarity error (CE): Smaller values of CE ensure that resources with a positive price maintain a balance between demand and supply, while resources priced at zero satisfy capacity constraints. We define
    $\textsf{CE}(\bar{\mathbf{x}},\mathbf{p})=\sqrt{\sum_{e\in\tilde{\mathcal{E}}}p_{e}^{2}z_{e}^{2}},$ (9)
    where ${z}_{e}=\sum_{u\in U}x_{u,e}-\ell_{e}$ is the excess demand. The definition of CE is motivated from the “complementarity condition” in general equilibrium theory in economics (Mas-Colell et al., 2006). We define $\textsf{tol}_{\textsf{CE}}$ to be the threshold for this error.
    $\textsf{ICE}(\bar{\mathbf{x}})={\max_{u\in U}\|\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}-1\|_{\infty}}.$ (10)
    We define $\textsf{tol}_{\textsf{ICE}}$ to be the threshold for this error.
  • Expected allocation error (EAE): Smaller value of EAE ensure that (4b) is satisfied. We define
    $\textsf{EAE}(\mathbf{x},\mathbf{y})=\max_{u\in U}\|\mathbf{y}_{u}-\mathbf{x}_{u}\|_{\infty}.$ (11)
    We define $\textsf{tol}_{\textsf{EAE}}$ to be the threshold for this error.
Remark 5.2.
Algorithm 1 bears resemblance to online learning algorithms for computing market equilibrium (Kleinberg and Leighton, 2003; Li et al., 2020; Bistritz and Bambos, 2021). In this process, AAM vehicles iteratively adjust their demand based on anonymized signals—such as prices, expected allocation, resource surplus, and dual multipliers—broadcasted by the SP. Meanwhile, the SP dynamically updates these signals in response to the aggregate demand from AAM vehicles, without requiring knowledge of their private valuations.
@@ 5.2 Step 2: Computing Integral Allocation
Figure 5: Flowchart of Algorithm 2, illustrating the ranking system and the removal of over-demanded edges.
Using the output from Algorithm 1, the service provider (SP) computes an integral allocation in a distributed manner using Algorithm 2. The SP sets the airspace price to $\mathbf{p}^{\dagger}$ (the output of Algorithm 1) and generates a ranked list of AAM vehicles based on $\bar{\mathbf{x}}^{\dagger}$ (cf. Sec. 5.2.1). This list is created by ranking the AAM vehicles in descending order according to the numerical value of the (fractional) allocation of their preferred routes (cf. Sec. 5.2.2).
5.2.1 Ranked List.
To generate a ranked list of AAM vehicles, we define $s^{\ast}(u)$ to be the most desired route of AAM vehicle $u$ in $R_{u}$. Using $\bar{\mathbf{x}}^{\dagger}$ from Algorithm 1, the SP creates a ranking over agents based on decreasing values of $x^{\dagger}_{u,e^{\ast}(s^{\ast}(u))}$, where $e^{\ast}(s^{\ast}(u))$ denotes the departing edge on the route $s^{\ast}(u)$. We denote the ranked list [^9] Ties are broken arbitrarily. of AAM vehicles by rank.
5.2.2 Integral Allocation.
After generating the ranked list, the SP fixes the prices for all resources based on the output of Algorithm 1 (i.e. $\bar{\mathbf{p}}^{\dagger}$ ) and iterates over AAM vehicles according to rank (cf. Line 3 in Algorithm 2). Each AAM vehicle is allocated its most desired feasible route (cf. Line 5-10 in Algorithm 2), subject to the current resource capacity. Suppose a capacity constraint is violated on any resource. In that case, the SP either removes that resource from the available pool for all remaining agents or increases its price to infinity for the remaining AAM vehicles (as described in line 12 of Algorithm 2). See Fig. 5 for a schematic illustration.
Algorithm 2 Integral Allocation Based on Ranked List
@@
2:Initialize remaining capacity: $\ell^{\text{rem}}\leftarrow\ell$
3:for $i=1$ to $|U|$ do
4:    $u\leftarrow\textsf{{rank}}_{i}$$\triangleright$ Select AAM vehicle from the ranked list
@@
7:         AAM vehicle $u$ reports its integral allocation $\mathbf{\bar{x}}_{u}$ by solving (2)
8:         if $x_{u,e}\leq\ell^{\text{rem}}_{e}$ for every $e\in\tilde{\mathcal{E}}$ then
9:             Update remaining capacity: $\ell^{\text{rem}}_{e}\leftarrow\ell^{\text{rem}}_{e}-\bar{x}_{u,e},\quad\forall e\in\tilde{\mathcal{E}}$
@@
11:         else
12:             Define contested set $C\leftarrow\{e\in\tilde{\mathcal{E}}\mid x_{u,e}>\ell^{\text{rem}}_{e}\}$triangleright Identify contested good
13:             for each $e\in C$ do
14:                 Update price: $p_{e}\leftarrow\infty$triangleright Remove $e$ from further consideration
15:             end for
16:         end if
17:    end while
18:end for
19:Return $\bar{\mathbf{x}}$
Remark 5.3.
For Algorithms 1-2, the SP only requires information on the resource demands of AAM vehicles, their feasible time trajectories, and their most preferred time trajectory. Importantly, the SP does not need any details regarding the private valuations of individual AAM vehicles. Moreover, each AAM vehicle does not have access to the demands or private valuations of other vehicles, ensuring that privacy is maintained throughout the process.
6 Drone Delivery in Toulouse: A Case Study
In this section, we validate our proposed mechanism using a dataset of simulated package deliveries by Airbus, as shown in Fig. 6. Particularly, we use a synthetic dataset generated by Airbus to simulate a drone-based package delivery scenario in Toulouse, France (Egorov et al., 2019; Chin et al., 2023a).
Dataset Specification: The data involves four warehouses located on the periphery of the city, which serve as hubs for UAV take-off and landing. Delivery requests are generated with spatial locations drawn uniformly across Toulouse and their temporal occurrences follow a Poisson process. Each request triggers a UAV to depart from the launchpad of a warehouse, deliver a package to the specified destination, and return to its origin. The UAV flight trajectory, including take-off, cruise, and landing phases, is generated by Airbus’s high-fidelity trajectory simulator, ensuring that the simulated operations closely mimic real-world conditions. The data set includes data corresponding to $177$ UAV flights, which spans roughly 6000 seconds ($100$ minutes). The average length of a UAV flight from a warehouse to the delivery location is approximately 300 seconds ($5$ minutes).
Airspace Specifications: To implement our auction approach, we partition Toulouse’s airspace into $12$ “cruising-altitude” regions, along with $4$ warehouses, as shown in Fig. 6. We construct a time-extended graph (cf. Definition 2.1) with a total of $T=400$ time-steps, where each step of corresponds to 15 seconds. Based on the data we find that the minimum capacity needed to accommodate all requests from UAVs is 14 units in all 12 cruising regions of the airspace and 4 units for vertiport departure and arrival at the warehouse locations. Therefore, to make this problem interesting, we set capacity to $50\%$ of this in all regions, unless otherwise specified.
UAV Specifications: Each UAV is either allocated a path on the time-extended graph or is “rebased.” For every UAV, the feasible set of paths on the time-extended graph includes its most desirable path, along with four alternative paths that each incur a one-time-step delay. If a UAV is rebased, then it converts its budget into outside options to be used again in the next auction window. Every rebased UAV requests access to the airspace from the start of the next auction window. Any UAV may be rebased at most twice, after which it is dropped and not considered in future auctions.
The private value generated by each UAV on its most desirable path is a uniformly random number between $150$ and $250$ units. The utility decreases by a factor of $0.95$ for each time-step delay in departure. If a UAV is rebased, its utility decreases by a factor of $0.5$. The utility derived by any UAV from dropping out is $40$ units.
Implementation Specifications: In each auction window, the SP allocates an additional budget of artificial currency to each participating UAV, randomly sampled between 150 and 250 units. This amount is then added to the UAV’s existing budget accumulated from past auctions. For our numerical study, we set the nominal tolerance values in Algorithm 1 as follows: $\textsf{tol}_{\textsf{CE}}$ is set to $0.1\%$ of its maximum possible value, $\textsf{tol}_{\textsf{ICE}}$ to $0.01\%$ of its maximum value, and $\textsf{tol}_{\textsf{EAE}}$ to $0.1\%$ of its maximum value. Additionally, we set $\beta=50$ and $N=30$ in Algorithm 1. We implemented our approach in Python and ran the simulations on a laptop equipped with a 12th-gen Intel Core i7-1200H CPU (14 cores, 20 threads) and 32GB DDR4 RAM. The operating system used was Ubuntu 22.04. Our code is publicly available at https://github.com/sastry-group/Mechanism-Design-for-AAM.
6.1 Qualitative Analysis
In this subsection, we present the outcome of our receding-horizon auction approach that we conducted for a total of $13$ (i.e, $I=13$) auctions. Before the start of every auction, the SP gathers the demand of AAM vehicles requesting access to the airspace. The overall budget of any UAV includes the new air credits they received and any unused air credits from a previous round. Additionally, we also compare our performance with two baselines.
In Fig. 7(a), we present the number of agents that were allocated, delayed, dropped, rebased-once and rebased-twice in different auctions. We see that the number of rebased and dropped agents increases as we proceed to later rounds. This is because the congestion builds up as we are operating in highly contested settings, as we set the maximum capacity of every region to be at $50\%.$ In Fig. 7(b), we present the market clearing error (MCE) after Algorithm 2 which captures the fraction of edges on the time-extended graphs which have positive price but for which the congestion is below the capacity. This metric aligns with the third component of the competitive equilibrium definition in Def. 4.1, capturing a common economic notion that we should not charge a payment on a good that is below its capacity. We see that this is a very small number, in the range of 0-0.6%, highlighting that our approach (Algorithm 1 + 2) is not imposing prices on the uncontested goods.
(a) Allocations for Each Auction
(b) Market Clearing Error After Integral Allocation (percent)
Figure 7: Properties of allocation finalized by our receding-horizon auction approach.
Table 1: Comparison to baseline auction approaches under different capacities
Approach Capacity Num. Times Rebased Num. Delayed UAVs Avg. Delay (time steps) Num. Rebased UAVs Avg. Times Rebased Num. Never Allocated
Budget-based 60% 20 12 1.83 17 1.18 0
50% 148 22 1.22 84 1.76 43
Profit-based 60% 59 17 2.41 32 1.84 24
50% 164 22 2.86 87 1.89 70
Ours 60% 12 10 2.3 12 1 0
50% 143 27 1.48 96 1.49 35
In Table 1, we compare our approach to two baselines based on the ascending clock auction (Cramton et al., 2006b). Both comparisons run simultaneous ascending clock auctions using $\beta$ as the price increment. Agents are allowed to perform price discovery by only bidding on their most beneficial request instead of all goods across their preferred and delayed requests. Due to (2b), we can assume agents will always bid on a request or the outside option and are therefore always active. The procedure is outlined in Algorithm 3. In the ”Budget-based” comparison, agents solve their IOP to determine their bid, and in the ”Profit-based” they determine their bid based on their profit (value - price). In each round, all agents bid, and prices are raised on contested goods until no more goods are contested.
Algorithm 3 Ascending Clock Auction Comparisons
1:Initialize prices: $\mathbf{p}\leftarrow\mathbf{0}$
2:repeat
3:    $\mathbf{\hat{x}}_{u}\leftarrow$ AAM vehicle $u$ integral output from (2) OR $\triangleright$ Budget-based Bid